When a number N is multiplied by another number which is 13 less than N, the result is 300. Find the number N.
Answer by Boreal(15235)

n*(n-13)=300
n^2-13n-300=0
(n-25)(n+12)=0
n=25 or -12
That will give 25 and 12=300 or -12 and -25, which also gives 300.